# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v2

@@ III-B Rejection-Sampled Universal Quantizer (RSUQ)
We will now review RSUQ [38], which is constructed based on the subtractive dithered quantizer (SDQ) [14, 57, 58].
Given a non-singular generator matrix $\mathbf{G}\in\mathbb{R}^{n\times n}$, a lattice is the set $\mathbf{G}\mathbb{Z}^{n}=\{\mathbf{G}\mathbf{j}:\,\mathbf{j}\in\mathbb{Z}^{n}\}$. A bounded set $\mathcal{P}\subseteq\mathbb{R}^{n}$ is called a basic cell of the lattice $\mathbf{G}\mathbb{Z}^{n}$ if $(\mathcal{P}+\mathbf{G}\mathbf{j})_{\mathbf{j}\in\mathbb{Z}^{n}}$ forms a partition of $\mathbb{R}^{n}$ [59, 60]. Specifically, the Voronoi cell $\mathcal{V}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\arg\min_{\mathbf{j}\in\mathbb{Z}^{n}}\|\mathbf{x}-\mathbf{G}\mathbf{j}\|=\mathbf{0}\}$ is a basic cell. Given a basic cell $\mathcal{P}$, we can define a lattice quantizer $Q_{\mathcal{P}}:\mathbb{R}^{n}\to\mathbf{G}\mathbb{Z}^{n}$ such that $Q_{\mathcal{P}}(\mathbf{x})=\mathbf{y}$ where $\mathbf{y}\in\mathbf{G}\mathbb{Z}^{n}$ is the unique lattice point that satisfies $\mathbf{x}\in-\mathcal{P}+\mathbf{y}$. The resulting quantization error $\mathbf{z}:=Q_{\mathcal{P}}(\mathbf{x})-\mathbf{x}$ depends deterministically on the input $\mathbf{x}$ and is approximately uniformly distributed over the basic cell of the lattice quantizer under some regularity assumptions. Therefore, it is often combined with probabilistic methods such as random dithering to construct SDQ [57, 58].
@@
A well-known property of SDQ is that the resulting quantization error is uniformly distributed over the basic cell of the quantizer and is statistically independent of the input signal [15, 58, 61, 62, 63].
@@
[38, Definition 4] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a subset $\mathcal{A}\subseteq\mathcal{P}$, and a sequence $S=(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}}$, $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ are i.i.d. dithers, the rejection-sampled universal quantizer (RSUQ) $Q_{\mathcal{A},\mathcal{P}}:\mathbb{R}^{n}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $\mathcal{A}$ against $\mathcal{P}$ is given by
@@
Note that SDQ is a special case of RSUQ where $\mathcal{A}=\mathcal{P}$. It is easy to check that the quantization error is uniform over $\mathcal{A}$ by using the standard rejection sampling argument and the “crypto” lemma [60, Lemma 4.1.1 and Theorem 4.1.1].
@@
[38, Proposition 5] For any $\mathbf{x}\in\mathbb{R}^{n}$, the quantization error $\mathbf{Z}:=Q_{\mathcal{A},\mathcal{P}}(\mathbf{x},S)-\mathbf{x}$ of RSUQ $Q_{\mathcal{A},\mathcal{P}}$, where $S\sim P_{S}$, follows the uniform distribution over the set $\mathcal{A}$, i.e., $\mathbf{Z}\sim\mathrm{Unif}(\mathcal{A})$.
By using layered construction as in [37, 36], RSUQ can be generalized to simulate an additive noise channel with noise following a continuous distribution. Consider a pdf $f:\mathbb{R}^{n}\to[0,\infty)$ and write its superlevel set as
@@
Let $f_{U}(u):=\mu(L_{u}^{+}(f))$ for $u>0$, which is also a pdf. If we generate $U\sim f_{U}$, and then $\mathbf{Z}|\{U=u\}\sim\mathrm{Unif}(L_{u}^{+}(f))$, then we have $\mathbf{Z}\sim f$ by the fundamental theorem of simulation [64].
@@
[38, Definition 12] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a probability density function $f:\mathbb{R}^{n}\to[0,\infty)$ where $L_{u}^{+}(f)$ is always bounded for $u>0$, and $\beta:(0,\infty)\to[0,\infty)$ satisfying $L_{u}^{+}(f)\subseteq\beta(u)\mathcal{P}$ for $u>0$, and a random pair $S=(U,(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}})$ where the latent variable $U\sim f_{U}$ with $f_{U}(u):=\mu(L_{u}^{+}(f))$, and $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ is a sequence of i.i.d. dither signals, the layered rejection-sampled universal quantizer (LRSUQ) $Q_{f,\mathcal{P}}:\mathbb{R}^{n}\times\mathbb{R}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $f$ against $\mathcal{P}$ is given by
@@
[38, Proposition 13] Consider LRSUQ $Q_{f,\mathcal{P}}$. For any random input $\mathbf{X}$, the quantization error defined by $\mathbf{Z}:=Q_{f,\mathcal{P}}(\mathbf{X},U,(\mathbf{V}_{i})_{i})-\mathbf{X}$, follows the pdf $f$, independent of $\mathbf{X}$.
@@ III-C Differential Privacy (DP)
Since clients do not directly transmit their data to the server, the original FL framework [1] provides a certain level of privacy. Nonetheless, a significant amount of information can still be inferred from the shared data, e.g., model parameters induced by gradient descent, by potential eavesdroppers within the FL network [65, 66, 67]. Consequently, a privacy mechanism, such as DP, is essential to protect the shared information.
In this section, we delve into a notion of (central) DP [10, 11]. [^5] Herein, DP refers to central DP. In DP, clients place trust in the server (or data curator) responsible for collecting and holding their individual data in a database $X\in\mathcal{X}$, where $\mathcal{X}$ denotes the collection of databases. The server then introduces privacy-preserving noise to the original datasets or query results through a randomized mechanism $\mathcal{F}$, producing an output $Y=\mathcal{F}(X)\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of possible outputs, before sharing them with untrusted data analysts. While this model requires a higher level of trust compared to the local model, it enables the design of significantly more accurate algorithms. Two databases $X$ and $X^{\prime}$ are considered adjacent if they differ in only one entry. More generally, we can define a symmetric adjacent relation $\mathcal{R}\subseteq\mathcal{X}^{2}$ and say that $X$ and $X^{\prime}$ are adjacent databases if $(X,X^{\prime})\in\mathcal{R}$. Here, we review the definition of $(\epsilon,\delta)$-differentially private, initially introduced by Dwork et al. [10].
Definition 6 ($(\epsilon,\delta)$-differential privacy [10]).
@@
As the applications of differential privacy continue to expand, the issue of managing the privacy budget is gaining increasing attention, particularly focusing on privacy composition and privacy amplification. Privacy composition [11, 68, 69] asserts that the privacy budgets of composition blocks accumulate, while privacy amplification [39, 40] provides tools to analyze and bound the privacy budget of a combination of a selected base privacy mechanism to be less than the privacy of its constituent parts.
In this work, we use the principle of privacy amplification by subsampling [39], whereby the privacy guarantees of a differentially private mechanism are amplified by applying it to a random subsample of the dataset. The problem of privacy amplification can be stated as follows. Let $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ be a privacy mechanism with privacy profile $\delta_{\mathcal{F}}$ with respect to the adjacent relation defined on $\mathcal{X}$, and let $s:\mathcal{W}\rightarrow\mathcal{X}$ be a subsampling mechanism. Consider the subsampled mechanism $\mathcal{F}^{s}:\mathcal{W}\rightarrow\mathcal{Y}$ given by $\mathcal{F}^{s}(X):=\mathcal{F}(s(X))$. The goal is to relate the privacy profile of $\mathcal{F}$ and $\mathcal{F}^{s}$.
For our purposes, it is instrumental to express differential privacy in terms of $e^{\epsilon}$-divergence [70]. Let $\mathfrak{M}(\mathcal{Y})$ denote the set of probability measures on the output space $Y$. The $e^{\epsilon}$-divergence between two probability measures $\mu,\mu^{\prime}\in\mathfrak{M}(\mathcal{Y})$ is defined as:
@@
To study the relavant properties of $\mathcal{F}$ from a privacy amplification point of view, we review the notions of privacy profile and group privacy profile [39] as follows. The privacy profile $\delta_{\mathcal{F}}$ of a mechanism $\mathcal{F}$ is a function associating to each privacy parameter $e^{\epsilon}$ a supremium on the $e^{\epsilon}$-divergence between the outputs of running the mechanism on two adjacent databases, i.e.,
@@
[…]
$\delta_{\mathcal{F},j}(\epsilon):=\sup_{d(X,X^{\prime})\leq k}D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime})),$ (11)
where $d(X,X^{\prime}):=\min\{k:\exists X_{1},\ldots,X_{k-1}\mathrm{\;with\;}(X,X_{1}),\ldots,(X_{k-1},X^{\prime})\in\mathcal{R}\}$. Note that $\delta_{\mathcal{F}}=\delta_{\mathcal{F},1}$.
We briefly review subsampling with replacement [39]. Let $\mathcal{U}$ be a set. We write $2^{\mathcal{U}}$ and $\mathbb{N}^{\mathcal{U}}$ for the collections of all sets and multisets, respectively, over $\mathcal{U}$. For an integer $n\geq 0$, we also write $2_{n}^{\mathcal{U}}$ and $\mathbb{N}_{n}^{\mathcal{U}}$ for the collections of all sets and multisets containing exactly $n$ elements, where the elements are counted with multiplicity for multisets. Given a multiset $X\in\mathbb{N}^{\mathcal{U}}$ we write $X_{a}$ for the number of occurrences of $a\in\mathcal{U}$ in $X$. The support of a multiset $X$ is the defined as the set $\mathrm{supp}(X):=\{a\in\mathcal{U}:X_{a}>0\}$. Given multisets $X,X^{\prime}\in\mathbb{N}^{\mathcal{U}}$ we write $X^{\prime}\subseteq X$ to denote that $X^{\prime}_{a}\leq X_{a}$ for all $a\in\mathcal{U}$. For order-independent datasets represented as multisets, the substitute-one relation, denoted as $X\simeq_{s}X^{\prime}$, holds whenever $\|X-X^{\prime}\|_{1}=2$ and $|X|=|X^{\prime}|$, i.e., $X^{\prime}$ is obtained by replacing an element in $X$ with a different element from $\mathcal{U}$. The subsampling with replacement mechanism $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ takes a set of $x$ of size $n$ and outputs a sample from the multinomial distribution $\omega=s(X)$ over all multisets $Y$ of size $\tau\leq n$ and $\mathrm{supp}(Y)\subseteq X$. We further assume that the base mechanism $\mathcal{F}:\mathbb{N}_{\tau}^{\mathcal{U}}\rightarrow\mathcal{Y}$ is defined on the multisets and has privacy profile $\delta_{\mathcal{F}}$ with respect to $\simeq_{s}$. In [39], Balle et al. demonstrated a bound on the privacy profile of the subsampled mechanism with replacement $\mathcal{F}^{s_{m}}:2_{n}^{\mathcal{U}}\rightarrow\mathcal{Y}$ with respect to $\simeq_{s}$ as follows.
@@
[39, Theorem 10] Let $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ be the subsampling with replacement mechanism and let $\mathcal{F}:{\color[rgb]{0,0,0}\mathbb{N}_{\tau}^{\mathcal{U}}}\rightarrow\mathcal{Y}$ be a base privacy mechanism. Then, for $\tilde{\epsilon}\geq 0$ and $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$, we have
@@
Recall that for a function $q:\mathcal{X}\to\mathbb{R}^{n}$ with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, the Gaussian mechanism $\mathcal{F}(X)=q(X)+\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ satisfies $(\epsilon,\delta)$-DP if $\tilde{\sigma}^{2}\geq 2\Delta_{2}^{2}\log(1.25/\delta)/\epsilon^{2}$ and $\epsilon\in(0,1)$ [11, Appendix A]. In [71], Balle and Wang gave a new analysis of the Gaussian mechanism that is valid for all $\epsilon>0$.
@@
[71, Theorem 8] Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the Gaussian mechanism $\mathcal{F}(X):=q(X)+\mathbf{Z}$ with $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, is $(\epsilon,\delta)$-DP if and only if
@@
Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, and let $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ be the subsampling with replacement mechanism. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the subsampled Gaussian mechanism ${\color[rgb]{0,0,0}\mathcal{F}^{s_{\tau}}(X)}=q\circ s(X)+N(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ is $(\epsilon,\delta)$-DP if and only if, for any $\tilde{\epsilon}>0$, $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$ and
@@
A standard group privacy analysis [72] gives the privacy profile

### v3

@@ III-B Rejection-Sampled Universal Quantizer (RSUQ)
We will now review RSUQ [Ling2024arxivRSUQ], which is constructed based on the subtractive dithered quantizer (SDQ) [roberts1962picture, ziv1985universal, zamir1992universal].
Given a non-singular generator matrix $\mathbf{G}\in\mathbb{R}^{n\times n}$, a lattice is the set $\mathbf{G}\mathbb{Z}^{n}=\{\mathbf{G}\mathbf{j}:\,\mathbf{j}\in\mathbb{Z}^{n}\}$. A bounded set $\mathcal{P}\subseteq\mathbb{R}^{n}$ is called a basic cell of the lattice $\mathbf{G}\mathbb{Z}^{n}$ if $(\mathcal{P}+\mathbf{G}\mathbf{j})_{\mathbf{j}\in\mathbb{Z}^{n}}$ forms a partition of $\mathbb{R}^{n}$ [conway2013sphere, zamir2014]. Specifically, the Voronoi cell $\mathcal{V}:=\{\mathbf{x}\in\mathbb{R}^{n}:\,\arg\min_{\mathbf{j}\in\mathbb{Z}^{n}}\|\mathbf{x}-\mathbf{G}\mathbf{j}\|=\mathbf{0}\}$ is a basic cell. Given a basic cell $\mathcal{P}$, we can define a lattice quantizer $Q_{\mathcal{P}}:\mathbb{R}^{n}\to\mathbf{G}\mathbb{Z}^{n}$ such that $Q_{\mathcal{P}}(\mathbf{x})=\mathbf{y}$ where $\mathbf{y}\in\mathbf{G}\mathbb{Z}^{n}$ is the unique lattice point that satisfies $\mathbf{x}\in-\mathcal{P}+\mathbf{y}$. The resulting quantization error $\mathbf{z}:=Q_{\mathcal{P}}(\mathbf{x})-\mathbf{x}$ depends deterministically on the input $\mathbf{x}$ and is approximately uniformly distributed over the basic cell of the lattice quantizer under some regularity assumptions. Therefore, it is often combined with probabilistic methods such as random dithering to construct SDQ [ziv1985universal, zamir1992universal].
@@
A well-known property of SDQ is that the resulting quantization error is uniformly distributed over the basic cell of the quantizer and is statistically independent of the input signal [schuchman1964dither, zamir1992universal, gray1993dithered, kirac1996results, zamir1996lattice].
@@
[Ling2024arxivRSUQ, Definition 4] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a subset $\mathcal{A}\subseteq\mathcal{P}$, and a sequence $S=(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}}$, $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ are i.i.d. dithers, the rejection-sampled universal quantizer (RSUQ) $Q_{\mathcal{A},\mathcal{P}}:\mathbb{R}^{n}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $\mathcal{A}$ against $\mathcal{P}$ is given by
@@
Note that SDQ is a special case of RSUQ where $\mathcal{A}=\mathcal{P}$. It is easy to check that the quantization error is uniform over $\mathcal{A}$ by using the standard rejection sampling argument and the “crypto” lemma [zamir2014, Lemma 4.1.1 and Theorem 4.1.1].
@@
[Ling2024arxivRSUQ, Proposition 5] For any $\mathbf{x}\in\mathbb{R}^{n}$, the quantization error $\mathbf{Z}:=Q_{\mathcal{A},\mathcal{P}}(\mathbf{x},S)-\mathbf{x}$ of RSUQ $Q_{\mathcal{A},\mathcal{P}}$, where $S\sim P_{S}$, follows the uniform distribution over the set $\mathcal{A}$, i.e., $\mathbf{Z}\sim\mathrm{Unif}(\mathcal{A})$.
By using layered construction as in [hegazy2022randomized, wilson2000layered], RSUQ can be generalized to simulate an additive noise channel with noise following a continuous distribution. Consider a pdf $f:\mathbb{R}^{n}\to[0,\infty)$ and write its superlevel set as
@@
Let $f_{U}(u):=\mu(L_{u}^{+}(f))$ for $u>0$, which is also a pdf. If we generate $U\sim f_{U}$, and then $\mathbf{Z}|\{U=u\}\sim\mathrm{Unif}(L_{u}^{+}(f))$, then we have $\mathbf{Z}\sim f$ by the fundamental theorem of simulation [robert2004monte].
@@
[Ling2024arxivRSUQ, Definition 12] Given a basic cell $\mathcal{P}$ of the lattice $\mathbf{G}\mathbb{Z}^{n}$, a probability density function $f:\mathbb{R}^{n}\to[0,\infty)$ where $L_{u}^{+}(f)$ is always bounded for $u>0$, and $\beta:(0,\infty)\to[0,\infty)$ satisfying $L_{u}^{+}(f)\subseteq\beta(u)\mathcal{P}$ for $u>0$, and a random pair $S=(U,(\mathbf{V}_{i})_{i\in\mathbb{N}^{+}})$ where the latent variable $U\sim f_{U}$ with $f_{U}(u):=\mu(L_{u}^{+}(f))$, and $\mathbf{V}_{1},\mathbf{V}_{2},\ldots\stackrel{{\scriptstyle iid}}{{\sim}}\mathrm{Unif}(\mathcal{P})$ is a sequence of i.i.d. dither signals, the layered rejection-sampled universal quantizer (LRSUQ) $Q_{f,\mathcal{P}}:\mathbb{R}^{n}\times\mathbb{R}\times\prod_{i\in\mathbb{N}^{+}}\mathcal{P}_{i}\to\mathbb{R}^{n}$ for $f$ against $\mathcal{P}$ is given by
@@
[Ling2024arxivRSUQ, Proposition 13] Consider LRSUQ $Q_{f,\mathcal{P}}$. For any random input $\mathbf{X}$, the quantization error defined by $\mathbf{Z}:=Q_{f,\mathcal{P}}(\mathbf{X},U,(\mathbf{V}_{i})_{i})-\mathbf{X}$, follows the pdf $f$, independent of $\mathbf{X}$.
@@ III-C Differential Privacy (DP)
Since clients do not directly transmit their data to the server, the original FL framework [McMahan2016FL] provides a certain level of privacy. Nonetheless, a significant amount of information can still be inferred from the shared data, e.g., model parameters induced by gradient descent, by potential eavesdroppers within the FL network [Zhu2019DeepLeak, zhao2020idlg, huang2021GradInvAtt]. Consequently, a privacy mechanism, such as DP, is essential to protect the shared information.
In this section, we delve into a notion of (central) DP [Dwork06DP, Dwork14Book]. [^5] Herein, DP refers to central DP. In DP, clients place trust in the server (or data curator) responsible for collecting and holding their individual data in a database $X\in\mathcal{X}$, where $\mathcal{X}$ denotes the collection of databases. The server then introduces privacy-preserving noise to the original datasets or query results through a randomized mechanism $\mathcal{F}$, producing an output $Y=\mathcal{F}(X)\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of possible outputs, before sharing them with untrusted data analysts. While this model requires a higher level of trust compared to the local model, it enables the design of significantly more accurate algorithms. Two databases $X$ and $X^{\prime}$ are considered adjacent if they differ in only one entry. More generally, we can define a symmetric adjacent relation $\mathcal{R}\subseteq\mathcal{X}^{2}$ and say that $X$ and $X^{\prime}$ are adjacent databases if $(X,X^{\prime})\in\mathcal{R}$. Here, we review the definition of $(\epsilon,\delta)$-differentially private, initially introduced by Dwork et al. [Dwork06DP].
Definition 6 ($(\epsilon,\delta)$-differential privacy [Dwork06DP]).
@@
As the applications of differential privacy continue to expand, the issue of managing the privacy budget is gaining increasing attention, particularly focusing on privacy composition and privacy amplification. Privacy composition [Dwork14Book, Abadi2016DL, Mironov2017RDP] asserts that the privacy budgets of composition blocks accumulate, while privacy amplification [Balle2018PrivAmpl, Feldman2018PrivAmpl] provides tools to analyze and bound the privacy budget of a combination of a selected base privacy mechanism to be less than the privacy of its constituent parts.
In this work, we use the principle of privacy amplification by subsampling [Balle2018PrivAmpl], whereby the privacy guarantees of a differentially private mechanism are amplified by applying it to a random subsample of the dataset. The problem of privacy amplification can be stated as follows. Let $\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ be a privacy mechanism with privacy profile $\delta_{\mathcal{F}}$ with respect to the adjacent relation defined on $\mathcal{X}$, and let $s:\mathcal{W}\rightarrow\mathcal{X}$ be a subsampling mechanism. Consider the subsampled mechanism $\mathcal{F}^{s}:\mathcal{W}\rightarrow\mathcal{Y}$ given by $\mathcal{F}^{s}(X):=\mathcal{F}(s(X))$. The goal is to relate the privacy profile of $\mathcal{F}$ and $\mathcal{F}^{s}$.
For our purposes, it is instrumental to express differential privacy in terms of $e^{\epsilon}$-divergence [Sason2016fdivInq]. Let $\mathfrak{M}(\mathcal{Y})$ denote the set of probability measures on the output space $Y$. The $e^{\epsilon}$-divergence between two probability measures $\mu,\mu^{\prime}\in\mathfrak{M}(\mathcal{Y})$ is defined as:
@@
To study the relavant properties of $\mathcal{F}$ from a privacy amplification point of view, we review the notions of privacy profile and group privacy profile [Balle2018PrivAmpl] as follows. The privacy profile $\delta_{\mathcal{F}}$ of a mechanism $\mathcal{F}$ is a function associating to each privacy parameter $e^{\epsilon}$ a supremium on the $e^{\epsilon}$-divergence between the outputs of running the mechanism on two adjacent databases, i.e.,
@@
The group privacy profile $\delta_{\mathcal{F},j}(\epsilon)$ ($j\geq 1$) is defined as:
$\delta_{\mathcal{F},j}(\epsilon):=\sup_{d(X,X^{\prime})\leq j}D_{e^{\epsilon}}(\mathcal{F}(X)||\mathcal{F}(X^{\prime})),$ (11)
where $d(X,X^{\prime}):=\min\{j:\exists X_{1},\ldots,X_{j-1}\mathrm{\;with\;}(X,X_{1}),\ldots,(X_{j-1},X^{\prime})\in\mathcal{R}\}$. Note that $\delta_{\mathcal{F}}=\delta_{\mathcal{F},1}$.
We briefly review subsampling with replacement [Balle2018PrivAmpl]. Let $\mathcal{U}$ be a set. We write $2^{\mathcal{U}}$ and $\mathbb{N}^{\mathcal{U}}$ for the collections of all sets and multisets, respectively, over $\mathcal{U}$. For an integer $n\geq 0$, we also write $2_{n}^{\mathcal{U}}$ and $\mathbb{N}_{n}^{\mathcal{U}}$ for the collections of all sets and multisets containing exactly $n$ elements, where the elements are counted with multiplicity for multisets. Given a multiset $X\in\mathbb{N}^{\mathcal{U}}$ we write $X_{a}$ for the number of occurrences of $a\in\mathcal{U}$ in $X$. The support of a multiset $X$ is the defined as the set $\mathrm{supp}(X):=\{a\in\mathcal{U}:X_{a}>0\}$. Given multisets $X,X^{\prime}\in\mathbb{N}^{\mathcal{U}}$ we write $X^{\prime}\subseteq X$ to denote that $X^{\prime}_{a}\leq X_{a}$ for all $a\in\mathcal{U}$. For order-independent datasets represented as multisets, the substitute-one relation, denoted as $X\simeq_{s}X^{\prime}$, holds whenever $\|X-X^{\prime}\|_{1}=2$ and $|X|=|X^{\prime}|$, i.e., $X^{\prime}$ is obtained by replacing an element in $X$ with a different element from $\mathcal{U}$. The subsampling with replacement mechanism $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ takes a set of $x$ of size $n$ and outputs a sample from the multinomial distribution $\omega=s(X)$ over all multisets $Y$ of size $\tau\leq n$ and $\mathrm{supp}(Y)\subseteq X$. We further assume that the base mechanism $\mathcal{F}:\mathbb{N}_{\tau}^{\mathcal{U}}\rightarrow\mathcal{Y}$ is defined on the multisets and has privacy profile $\delta_{\mathcal{F}}$ with respect to $\simeq_{s}$. In [Balle2018PrivAmpl], Balle et al. demonstrated a bound on the privacy profile of the subsampled mechanism with replacement $\mathcal{F}^{s_{m}}:2_{n}^{\mathcal{U}}\rightarrow\mathcal{Y}$ with respect to $\simeq_{s}$ as follows.
@@
[Balle2018PrivAmpl, Theorem 10] Let $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ be the subsampling with replacement mechanism and let $\mathcal{F}:\mathbb{N}_{\tau}^{\mathcal{U}}\rightarrow\mathcal{Y}$ be a base privacy mechanism. Then, for $\tilde{\epsilon}\geq 0$ and $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$, we have
@@
Recall that for a function $q:\mathcal{X}\to\mathbb{R}^{n}$ with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, the Gaussian mechanism $\mathcal{F}(X)=q(X)+\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ satisfies $(\epsilon,\delta)$-DP if $\tilde{\sigma}^{2}\geq 2\Delta_{2}^{2}\log(1.25/\delta)/\epsilon^{2}$ and $\epsilon\in(0,1)$ [Dwork14Book, Appendix A]. In [balle2018ImprovGauss], Balle and Wang gave a new analysis of the Gaussian mechanism that is valid for all $\epsilon>0$.
@@
[balle2018ImprovGauss, Theorem 8] Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the Gaussian mechanism $\mathcal{F}(X):=q(X)+\mathbf{Z}$ with $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, is $(\epsilon,\delta)$-DP if and only if
@@
Let $q:\mathcal{X}\rightarrow\mathbb{R}^{n}$ be a function with global $\ell_{2}$-sensitivity $\Delta_{2}:=\sup_{(X,X^{\prime})\in\mathcal{R}}\|q(X)-q(X^{\prime})\|_{2}$, and let $s_{\tau}:2_{n}^{\mathcal{U}}\rightarrow\mathfrak{M}(\mathbb{N}_{\tau}^{\mathcal{U}})$ be the subsampling with replacement mechanism. Then, for any $\epsilon\geq 0$ and $\delta\in[0,1]$, the subsampled Gaussian mechanism $\mathcal{F}^{s_{\tau}}(X)=q\circ s(X)+N(\mathbf{0},\tilde{\sigma}^{2}\mathbf{I}_{n})$ is $(\epsilon,\delta)$-DP if and only if, for any $\tilde{\epsilon}>0$, $\epsilon=\log\left(1+p(e^{\tilde{\epsilon}}-1)\right)$ where $p:=1-\left(1-\frac{1}{n}\right)^{\tau}$ and
@@
A standard group privacy analysis [Vadhan2017] gives the privacy profile